OpenStudy (anonymous):

Notice of a town meeting is spread by word of mouth to the town's 4000 residents. After t days, the number of people who know about the meeting is expressed by the function f(t) = 4000/(1 + 40 e^(-0.4 t)) How long will it take for half of the townspeople to have heard about the meeting?

OpenStudy (anonymous):

right now im at: f(t) = 4000/(1 + 40 e^(-0.4 t)).

OpenStudy (anonymous):

do i plug in something for t then solve?

OpenStudy (anonymous):

2, 000 = 4000/(1 + 40 e^(-0.4 t))

OpenStudy (anonymous):

yeah since it says half of pop but what then?

OpenStudy (anonymous):

Calculate!

OpenStudy (anonymous):

f(t) = 2000/(1 + 40 e^(-0.4 t)).

OpenStudy (anonymous):

1 = 2/[ 1 + 40 e^(-0.4 t) ]

OpenStudy (anonymous):

arent i supposed to plug in a number for t?

OpenStudy (anonymous):

Did you see I plug 2,000 for f(t) which number of half of town?

OpenStudy (anonymous):

Then find t?

OpenStudy (anonymous):

yeah but i dont plug that in 2000 for t though right?

OpenStudy (anonymous):

Of cousrse not, because 2,000 is number of person f (t), not time t!

OpenStudy (anonymous):

Let me know your result!

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

Firstly, do you know how to plug in ? Pls show me!